Find the mass of nonelectrolyte solute dissolved in 315 g water with boiling point temperature of 100.45 ⁰C. The molar mass of solute is 46 g/mole

                          Kb = 0.52 ⁰C/m  and Tb = 100 ⁰C (boiling point of pure solvent)




1
Expert's answer
2022-02-06T11:20:27-0500

Delta TB = i × kb × m .

( 100.45 - 100 ) = 1 × 0.52 × ( y / 46 ) / 0.315


0.45 = 1 × 0.52 × y / 46 × 0.315


y = ( 0.45 × 46 × 0.315 / 0.52 ) g .


y = ( 12.53 ) g